OpenStudy (anonymous):

(x+5)/(x-2)=(x-2)/(x+4)... last one

OpenStudy (saifoo.khan):

simply cross multiply.

OpenStudy (saifoo.khan):

(x+5) (x+4) ------ * ------ (x-2) (x-2)

OpenStudy (anonymous):

so just multiply x+5 times x-2? and the other way around?

OpenStudy (saifoo.khan):

now just multiply.. (x+5)(x+4) ---------- (x-2)(x-2)

OpenStudy (anonymous):

still don't get it lol

OpenStudy (saifoo.khan):

Lol, i simply cross multiplied. write this down in our book, then u will understand.

jhonyy9 (jhonyy9):

solve this (x+5)(x+4)=(x-2)(x-2)

OpenStudy (saifoo.khan):

yea, u can do this too.

jhonyy9 (jhonyy9):

x^2 +9x +20 =x^2 +2x -8

jhonyy9 (jhonyy9):

7x=-28

jhonyy9 (jhonyy9):

x=-28/7 = -4 x=-4

OpenStudy (anonymous):

ok I see where you got the 20 and the 8 but where did the 9 come from?

OpenStudy (saifoo.khan):

after multiplying them. FOIL.

jhonyy9 (jhonyy9):

but x not can being -4 because than the fraction (x-2)/x-4) is undefined

jhonyy9 (jhonyy9):

sorry the fraction x-2/x+4

jhonyy9 (jhonyy9):

because than the denominator will be zero

OpenStudy (anonymous):

OHHH ok I remember this now.